### v6

# Dominated Actions in Imperfect-Information Games

Abstract
Dominance is a fundamental concept in game theory. In normal-form games dominated strategies can be identified in polynomial time. As a consequence, iterative removal of dominated strategies can be performed efficiently as a preprocessing step for reducing the size of a game before computing a Nash equilibrium. For imperfect-information games in extensive form, we could convert the game to normal form and then iteratively remove dominated strategies in the same way; however, this conversion may cause an exponential blowup in game size. In this paper we define and study the concept of dominated actions in imperfect-information games. Our main result is a polynomial-time algorithm for determining whether an action is dominated (strictly or weakly) by any mixed strategy in $n$-player games, which can be extended to an algorithm for iteratively removing dominated actions. This allows us to efficiently reduce the size of the game tree as a preprocessing step for Nash equilibrium computation. We explore the role of dominated actions empirically in “All In or Fold” No-Limit Texas Hold’em poker.
1 Introduction
In a normal-form game, (mixed) strategy $\sigma_{i}$ for player $i$ is strictly dominated if there exists another (mixed) strategy $\sigma^{\prime}_{i}$ such that $\sigma^{\prime}_{i}$ performs strictly better than $\sigma_{i}$ regardless of the strategies used by the opponent(s): formally, if $u_{i}(\sigma^{\prime}_{i},s_{-i})>u_{i}(\sigma_{i},s_{-i)}$ for all pure strategy profiles (vectors of pure strategies) $s_{-i}\in S_{-i}$ for the opposing agents. (Note that the requirement that this holds for all opposing pure strategy profiles is enough to ensure that it also holds for all mixed strategy profiles of the opponents as well.) Clearly it would be irrational for an agent to play a strictly dominated strategy, as another strategy will do strictly better regardless of the beliefs of what the opponents would play. Strategy $\sigma^{\prime}_{i}$ weakly dominates $\sigma$ if the inequality holds weakly (though strictly for at least one pure strategy profile): formally, $u_{i}(\sigma^{\prime},s_{-i})\geq u_{i}(\sigma_{i},s_{-i})$ for all $s_{-i}\in S_{-i}$ where the inequality is strict for at least one $s_{-i}$. (The condition that at least one inequality is strict is simply to rule out saying that a strategy is dominated by an identical strategy).
It seems natural to simplify a game by eliminating strategies that are dominated from the game to reduce its size and focus analysis on a smaller game. It can easily be shown that applying an iterative process of removing one dominated strategy for one player, then removing one for another (or possibly the same) player in the reduced game, etc., will ultimately result in a smaller game that contains a Nash equilibrium from the original game. For this reason, this procedure of iterated removal of dominated strategies is often performed as a preprocessing step to reduce the size of a game before computing a Nash equilibrium (or other desired solution concept). As it turns out, all processes of iterated removal of strictly dominated strategies produce the same reduced game regardless of the order of elimination in finite games (while this is not necessarily the case for iterated removal of weakly dominated strategies) [6, 9]. While iterated removal of weakly dominated strategies can sometimes reduce the number of equilibria, it can never create new equilibria, and therefore even this procedure is very useful as a preprocessing step for Nash equilibrium computation.
There is a linear-time algorithm for determining whether a (mixed) strategy $\sigma_{i}$ is strictly dominated by any pure strategy for player $i$ [14]. This algorithm simply iterates over each pure strategy $s_{i}$ for player $i$ and tests whether it performs strictly better than $\sigma_{i}$ against each opposing pure strategy profile $s_{-i}$. This procedure has complexity $O(|A|)$, where $A=\times_{i}A_{i}$ is the set of joint pure strategy profiles, and so takes time linear in the size of the game. The procedure can also be easily adapted to produce an algorithm for determining whether a mixed strategy profile is weakly dominated by any pure strategy in linear time. Note that it is possible for a strategy to be dominated by a mixed strategy and not be dominated by any pure strategy [14]. In order to determine whether a (mixed) strategy is strictly dominated by a mixed strategy, while the above procedure does not work, it turns out that there exists a linear programming formulation that runs in polynomial time, and there also exists a linear programming formulation that determines whether a (mixed) strategy is weakly dominated by a mixed strategy [1]. Therefore, regardless of whether the strategies being tested as dominated or dominating are mixed or pure, it can be checked in polynomial time. [^1] However, most other computational questions related to dominance can be solved in polynomial time for strict but are NP-hard for weak dominance. For example, determining whether there exists some elimination path under which strategy $s_{i}$ is eliminated, given strategy subsets $S^{\prime}_{i}\subset S_{i}$ determining whether there exists a maximally reduced game where each player has the pure strategies $S^{\prime}_{i}$, and given constants $k_{i}$ determining whether there exists a maximally reduced game where each player has exactly $k_{i}$ pure strategies [14, 7].
Dominance is well-understood in normal-form games, where classical characterizations such as those of Pearce [12] describe exactly when a pure strategy is strictly or weakly dominated by a mixed strategy of the opponent. These characterizations, however, do not extend to extensive-form games. In imperfect-information settings, the dominance of an action depends on which opponent strategies reach the corresponding information set, a feature with no analogue in the normal-form framework. Consequently, an action in an extensive-form game may be locally dominated even when every pure or mixed strategy containing that action is undominated in the induced normal form. This distinction highlights the need for dominance notions and algorithms tailored specifically to extensive-form games.
2 Extensive-form games
While the normal form can be used to model simultaneous moves, another representation, called the extensive form, is generally preferred when modeling settings that have sequential moves. The extensive form can also model simultaneous moves, as well as chance events and imperfect information (i.e., situations where some information is available to only some of the agents and not to others). Extensive-form games consist primarily of a game tree; each non-terminal node has an associated player (possibly chance) that makes the decision at that node, and each terminal node has associated utilities for the players. Additionally, game states are partitioned into information sets, where the player whose turn it is to move cannot distinguish among the states in the same information set. Therefore, in any given information set, a player must choose actions with the same distribution at each state contained in the information set. If no player forgets information that they previously knew, we say that the game has perfect recall. A (behavioral) strategy for player $i,$ $\sigma_{i}\in\Sigma_{i},$ is a function that assigns a probability distribution over all actions at each information set belonging to $i$. In theory, every extensive-form game can be converted to an equivalent normal-form game; however, there is an exponential blowup in the size of the game representation, and therefore such a conversion is undesirable. Instead, new algorithms have been developed that operate on the extensive form representation directly. It turns out that the complexity of computing equilibria in extensive-form games is similar to that of normal-form games; a Nash equilibrium can be computed in polynomial time in two-player zero-sum games (with perfect recall) [8], while the problem is hard for two-player general-sum and multiplayer games. One algorithm for computing an equilibrium in two-player zero-sum extensive-form games with perfect recall is based on solving a linear programming formulation [8]. This formulation works by modeling each sequence of actions for each player as a variable, and is often called the sequence form LP algorithm. Note that while the number of pure strategies is exponential in the size of the game tree, the number of action sequences is linear. The method uses several matrices defined as follows. For player 1, the matrix $\mathbf{E}$ is defined where each row corresponds to an information set (including an initial row for the “empty” information set), and each column corresponds to an action sequence (including an initial row for the “empty” action sequence). In the first row of $\mathbf{E}$ the first element is 1 and all other elements are 0; subsequent rows have -1 for the entries corresponding to the action sequence leading to the root of the information set, and 1 for all actions that can be taken at the information set (and 0 otherwise). Thus $\mathbf{E}$ has dimension $c_{1}\times d_{1}$, where $c_{i}$ is the number of information sets for player $i$ and $d_{i}$ is the number of action sequences for player $i$. Matrix $\mathbf{F}$ is defined analogously for player 2. The vector $\mathbf{e}$ is defined to be a column vector of length $c_{1}$ with 1 in the first position and 0 in other entries, and vector $\mathbf{f}$ is defined with length $c_{2}$ analogously. The matrix $\mathbf{A}$ is defined with dimension $d_{1}\times d_{2}$ where entry $A_{ij}$ gives the payoff for player 1 when player 1 plays action sequence $i$ and player 2 plays action sequence $j$ multiplied by the probabilities of chance moves along the path of play. The matrix $\mathbf{B}$ of player 2’s expected payoffs is defined analogously. In zero-sum games $\mathbf{B}=-\mathbf{A}.$
Given these matrices we can solve one of two linear programming problems to compute a Nash equilibrium in zero-sum extensive-form games [8]. In the first formulation the primal variables $\mathbf{x}$ correspond to player 1’s mixed strategy while the dual variables correspond to player 2’s strategy. In the second formulation, which is the dual problem of the first formulation, the primal decision variables $\mathbf{y}$ correspond to player 2’s strategy while the dual variables correspond to player 1’s strategy.
[…]
3 Dominated actions
In extensive-form games, we can consider analogous concepts of strict, weak, and iterated dominance of strategies as for normal-form games. However, unlike in the normal-form setting, identification of a dominated extensive-form strategy does not necessarily allow us to reduce the size of the game, since it is possible that some of the actions played by the dominated strategy are also played by non-dominated strategies. In order to obtain the computational advantage of game size reduction, we must consider a stronger concept of dominated actions. We first present several plausible candidate definitions for dominated actions which we demonstrate to be problematic. Our first candidate definition is given as Candidate Definition 1. This definition states that action $a_{i}$ for player $i$ at information set $I_{i}$ is strictly dominated by action $b_{i}$ at the same information set if every leaf node succeeding $a_{i}$ produces a strictly smaller payoff for player $i$ than every leaf node succeeding action $b_{i}.$ An analogous definition for weak dominance is given in Candidate Definition 2.
Candidate Definition 1.
If for every leaf node $N^{a_{i}}$ that follows action $a_{i}$ for player $i$ at information set $I_{i}$ and every leaf node $N^{b_{i}}$ that follows action $b_{i}$ for player $i$ at the same information set $I_{i}$, $u_{i}\left(N^{b_{i}}\right)>u_{i}\left(N^{a_{i}}\right),$ then $b_{i}$ strictly dominates $a_{i}.$
Candidate Definition 2.
If for every leaf node $N^{a_{i}}$ that follows action $a_{i}$ for player $i$ at information set $I_{i}$ and every leaf node $N^{b_{i}}$ that follows action $b_{i}$ for player $i$ at the same information set $I_{i}$, $u_{i}\left(N^{b_{i}}\right)\geq u_{i}\left(N^{a_{i}}\right)$ where inequality is strict for at least one node, then $b_{i}$ weakly dominates $a_{i}.$
The problem with these definitions is that they are too strong; it is still possible for player 1 to strictly prefer action $b_{i}$ to $a_{i}$ regardless of the strategy used by player 2 even if Candidate Definition 1 does not hold. This is illustrated in the following game depicted in Figure 1. In this game, chance makes an initial move taking each of two actions with probability $\frac{1}{2}.$ Then player 1 (red) selects one of two actions at a single information set. Player 2 (blue) then takes one of two actions after observing both the moves of chance and player 1. It is clear that action 2 for player 2 at their top information set and action 1 for player 2 at their bottom information set are both strictly dominated according to Candidate Definition 1. The smaller game obtained after removing these actions is depicted in Figure 2. In the smaller game, the expected utility of playing action 1 is 0.5(100)+0.5(100)=00.5(-100)+0.5(100)=0 and the expected utility of playing action 2 is 0.5(50)+0.5(50)=50.0.5(-50)+0.5(-50)=-50. Since player 2 does not take any actions, player 1 always achieves a strictly higher expected utility by playing action 1. However, action 2 is not strictly dominated according to Candidate Definition 1 because in one leaf node succeeding action 1 player 1 obtains payoff -100 which is lower than the payoff of -50 at leaf nodes following action 2.
Figure 1: Example two-player imperfect-information extensive-form game.
Figure 2: Result of removing two dominated actions from game in Figure 1.
Candidate Definitions 1 and 2 provide sufficient conditions for an action to be dominated, but we have demonstrated that it is clearly possible for actions that do not satisfy these definitions to also be dominated. Thus these conditions are too strong, and we will refer to strategies that satisfy them as being strongly dominated (strictly or weakly, respectively). The concept of strong dominance is not without merit, as it can be verified very efficiently by simply iterating over the leaf nodes succeeding each action. Thus it may be useful to first remove actions that are strongly dominated as a preprocessing step before performing potentially more costly computations to remove other actions.
We next consider candidate definition given by Candidate Definition 3, where $u_{i}$ denotes the expected utility accounting for randomized moves of chance as well as potential randomization in the players’ strategies. This definition states that action $a_{i}$ for player $i$ at information set $I_{i}$ is strictly dominated (potentially by a probability distribution over other actions at the same information set), if there exists a strategy $\sigma^{-a_{i}}_{i}$ that never plays $a_{i}$ at $I_{i}$ that always has strictly higher expected utility than every strategy that plays $a_{i}$ at $I_{i}.$ Again this definition clearly provides a sufficient condition for $a_{i}$ to be dominated; however, the issue with this definition is that the strategies may potentially take actions early in the game tree that prevent the game from ever reaching information set $I_{i}.$ Consider the simple example game in Figure 3. Suppose we want to apply Candidate Definition 3 to determine whether action 2 is strictly dominated for player 2. Then $\sigma^{-a_{i}}_{i}$ is the strategy for player 2 that plays action 1 with probability 1, and $\sigma^{a_{i}}_{i}$ must be the strategy for player 2 that plays action 2 with probability 1. Now suppose that player 1 plays action 2 with probability 1, which we denote as $\sigma_{-i}$. Then clearly $u_{i}\left(\sigma^{-a_{i}}_{i},\sigma_{-i}\right)=u_{i}\left(\sigma^{a_{i}}_{i},\sigma_{-i}\right)=0,$ since both strategy profiles will result in reaching the bottom leaf node yielding payoff 0.
Candidate Definition 3.
Action $a_{i}$ for player $i$ is strictly dominated at information set $I_{i}$ if there exists a mixed strategy $\sigma^{-a_{i}}_{i}$ that plays action $a_{i}$ at $I_{i}$ with probability 0 such that for every mixed strategy $\sigma^{a_{i}}_{i}$ for player $i$ that plays action $a_{i}$ with probability 1 at $I_{i}$, $u_{i}\left(\sigma^{-a_{i}}_{i},\sigma_{-i}\right)>u_{i}\left(\sigma^{a_{i}}_{i},\sigma_{-i}\right)$ for all opposing strategies $\sigma_{-i}\in\Sigma_{-i}.$
Figure 3: Extensive-form game demonstrating problem with Candidate Definition 3.
The problem with Candidate Definition 3 is that it allows the players to deviate from the path of play leading to the relevant information set $I_{i}.$ We address this limitation in our new definitions. Strictly-dominated actions are defined in Definition 1 and weakly-dominated actions are defined in Definition 2. In these definitions $\Sigma^{I_{i}}_{-i}$ denotes the set of mixed strategy profiles for the opponents that always take actions leading to information set $I_{i}$ when possible. Note that these definitions apply to games with any number of players. They also consider actions that are dominated by any mixed strategy (not necessarily just a pure action at $I_{i}$). It is easy to verify that these definitions address the issues that arose in the examples from Figure 1 and Figure 3. We can apply these definitions repeatedly in succession to perform iterative removal of dominated actions, as for the normal form.
Definition 1.
Action $a_{i}$ for player $i$ is strictly dominated at information set $I_{i}$ if there exists a mixed strategy $\sigma^{-a_{i}}_{i}$ that always plays to get to $I_{i}$ and plays action $a_{i}$ at $I_{i}$ with probability 0 such that for every mixed strategy $\sigma^{a_{i}}_{i}$ for player $i$ that always plays to get to $I_{i}$ and plays action $a_{i}$ with probability 1 at $I_{i}$, $u_{i}\left(\sigma^{-a_{i}}_{i},\sigma_{-i}\right)>u_{i}\left(\sigma^{a_{i}}_{i},\sigma_{-i}\right)$ for all opposing strategies $\sigma_{-i}\in\Sigma^{I_{i}}_{-i}.$
Definition 2.
Action $a_{i}$ for player $i$ is weakly dominated at information set $I_{i}$ if there exists a mixed strategy $\sigma^{-a_{i}}_{i}$ that always plays to get to $I_{i}$ and plays action $a_{i}$ at $I_{i}$ with probability 0 such that for every mixed strategy $\sigma^{a_{i}}_{i}$ for player $i$ that always plays to get to $I_{i}$ and plays action $a_{i}$ with probability 1 at $I_{i}$, $u_{i}\left(\sigma^{-a_{i}}_{i},\sigma_{-i}\right)\geq u_{i}\left(\sigma^{a_{i}}_{i},\sigma_{-i}\right)$ for all opposing strategies $\sigma_{-i}\in\Sigma^{I_{i}}_{-i}$, where the inequality is strict for at least one $\sigma_{-i}.$
The example games considered were created using the open-source software package Gambit [13]. Gambit has tools that allow the user to remove “strictly dominated” or “strictly or weakly dominated” actions, and these procedures can be repeated to iteratively remove multiple actions sequentially; however, there is no documentation regarding the algorithms applied or definitions of dominance used. In the example game from Figure 1, Gambit correctly identifies that action 2 for player 2 at their top information set and action 1 for player 2 at their bottom information set are both strictly dominated, and removes these to construct the smaller game in Figure 2. However, as for strong domination, Gambit fails to recognize that action 2 is strictly dominated by action 1 for player 1 in the reduced game. This example demonstrates that Gambit’s procedure does not remove all dominated actions (though it does not necessarily imply that Gambit is only removing strongly dominated actions).
4 Algorithm for identifying dominated actions in two-player games
Suppose we want to determine whether an action $c$ for player 1 is dominated at information set $I$ in a two-player perfect-recall extensive-form game $G$. Using the sequence form representation, suppose that action $c$ is the final action in the sequence with index $i$ for player 1. Let $S_{I}$ denote the set of indices corresponding to action sequences leading to $I$. We would like to solve the following problem.
[…]
Consider the following problem, where now player 2 controls two action sequences $\mathbf{y_{1}},\mathbf{y_{2}}$:
[…]
Proposition 1.
The optimal objective values in Problem 1 and Problem 2 are the same.
Proof.
Let $f_{1}$ be the optimal objective value in Problem 1 and $f_{2}$ be the optimal objective value in Problem 2. Suppose that the optimal variables in Problem 1 are $\mathbf{x}^{1}_{1},\mathbf{x}^{1}_{2},\mathbf{y}^{1}.$ Now set $\mathbf{x}^{2}_{1}=\mathbf{x}^{1}_{1},$ $\mathbf{x}^{2}_{2}=\mathbf{x}^{1}_{2},$ $\mathbf{y}^{2}_{1}=\mathbf{y}^{1},$ $\mathbf{y}^{2}_{2}=\mathbf{y}^{1}.$ Then $(\mathbf{x}^{2}_{1},\mathbf{x}^{2}_{2},\mathbf{y}^{2}_{1},\mathbf{y}^{2}_{2})$ gives a feasible solution to Problem 2 with objective value equal to $f_{1}.$ So $f_{2}\geq f_{1}$. Now suppose that the optimal variables in Problem 2 are $\mathbf{x}^{2}_{1},\mathbf{x}^{2}_{2},\mathbf{y}^{2}_{1},\mathbf{y}^{2}_{2}.$ Now set $\mathbf{x}^{1}_{1}=\mathbf{x}^{2}_{1},$ $\mathbf{x}^{1}_{2}=\mathbf{x}^{2}_{2},$ and set $\mathbf{y}^{1}$ equal to the strategy that follows $\mathbf{y}^{2}_{1}$ at states following action $c$ for player 1, and follows $\mathbf{y}^{2}_{2}$ otherwise. Then $(\mathbf{x}^{1}_{1})^{T}\mathbf{A}\mathbf{y}^{1}=(\mathbf{x}^{1}_{1})^{T}\mathbf{A}\mathbf{y}^{2}_{1},$ since both players only take actions to get to information set $I$, $\mathbf{x}^{1}_{1}$ takes action $c$ at $I$, and the strategies $\mathbf{y}^{1}$ and $\mathbf{y}^{2}_{1}$ are identical after player 1 takes action $c$ at $I.$ Similarly, $(\mathbf{x}^{1}_{2})^{T}\mathbf{A}\mathbf{y}^{1}=(\mathbf{x}^{1}_{2})^{T}\mathbf{A}\mathbf{y}^{2}_{2},$ since both players only take actions to get to $I$, $\mathbf{x}^{1}_{1}$ does not take action $c$ at $I$, and the strategies $\mathbf{y}^{1}$ and $\mathbf{y}^{2}_{2}$ are identical after player 1 does not take action $c$ at $I.$ So $f_{1}\geq f_{2}$. So we conclude that $f_{1}=f_{2}.$
Proposition 1 allows us to divide Problem 2 into the following two subproblems. If $f$ is the optimal objective value of Problem 2, $f_{1}$ is the optimal objective of Problem 3, and $f_{2}$ is the optimal objective of Problem 4, then we have $f=f_{1}-f_{2}.$
[…]
Let us first look at Problem 3 and consider the inner subproblem for a fixed $\mathbf{x}_{2}$.
[…]
The Lagrangian is
$L(\mathbf{y},\boldsymbol{\lambda},\boldsymbol{\gamma},\mathbf{r})=\mathbf{x}^{T}_{2}\mathbf{A}\mathbf{y}-(\mathbf{f}^{T}-\mathbf{y}^{T}\mathbf{F}^{T})\boldsymbol{\lambda}-\sum_{k\in S_{I}}\boldsymbol{\gamma}_{k}(\mathbf{y}_{k}-1)-\mathbf{r}^{T}\mathbf{y}$
$\frac{\partial L}{\partial\mathbf{y}}=\mathbf{x}^{T}_{2}\mathbf{A}+\boldsymbol{\lambda}^{T}\mathbf{F}-\sum_{k\in S_{I}}\boldsymbol{\gamma}_{k}\mathbf{e_{k}}-\mathbf{r}^{T}$
The dual problem is
[…]
So Problem 3 is equivalent to:
[…]
Next consider Problem 4. Note that both players are aligned in trying to maximize the objective. Let us define a new problem $\overline{G}$ where player 1 selects the actions for both players. We denote player 1’s strategy in this modified game by $\overline{\mathbf{x}}_{1}.$ We modify $\mathbf{E}$ and resize $\mathbf{e}$ accordingly and denote them by $\mathbf{E^{\prime}}$ and $\mathbf{e^{\prime}}.$ The payoffs can now be represented as a vector $\mathbf{a}$. In the new representation let $i^{\prime}$ denote the index of the sequence with concluding action $c$, and let $S^{\prime}_{I}$ denote the set of indices corresponding to action sequences leading to $I$.
[…]
Let $u_{2}$ denote the optimal objective value for optimization problem 5, and $u_{1}$ denote the optimal objective value for problem 6. If $u_{2}>u_{1}$ then we conclude that action $c$ is strictly dominated. If $u_{2}<u_{1}$ then we conclude that action $c$ is not strictly or weakly dominated. If $u_{2}=u_{1}$, let $u_{3}$ denote the optimal objective value for problem 7, and let $u_{4}$ denote the optimal objective value for problem 8. If $u_{3}>u_{4}$ then we conclude that action $c$ is weakly dominated, and if $u_{3}=u_{4}$ we conclude that action $c$ is not strictly or weakly dominated (note that we cannot have $u_{3}<u_{4}$).
[…]
We can perform this procedure for every action at each information set of player 1, and analogously for player 2. Since the number of actions is linear in the size of the game tree, the overall procedure involves solving a linear number of linear programs and therefore runs in polynomial time. We can repeat the procedure iteratively until no more actions are removed for any player. Thus, iterative removal of dominated actions can be performed in polynomial time. Note that the procedure applies to all two-player games and does not assume that they are zero sum. The procedure also removes actions that are dominated by any mixed strategy (which may play a probability distribution over actions at the same information set $I$), not just actions that are dominated by a pure action.
Theorem 1.
There exists a polynomial-time algorithm that determines whether a given action is strictly dominated in a two-player perfect-recall extensive-form game of imperfect information.
Theorem 2.
There exists a polynomial-time algorithm that determines whether a given action is weakly dominated in a two-player perfect-recall extensive-form game of imperfect information.
Theorem 3.
Iterated removal of strictly and weakly dominated actions can be performed in polynomial time in a two-player perfect-recall extensive-form game of imperfect information.
5 Extension to $n$ players
Suppose we wish to determine whether an action of player 1 is strictly or weakly dominated in an $n$-player perfect-recall game $\Gamma$ with $n>2$. Although the linear-programming characterization developed above applies only to two-player games with perfect recall, we can reduce the general case to the two-player setting by merging all players other than player 1 into a single aggregated opponent. This merged player controls every decision node that originally belonged to players $2,\dots,n$, and its information sets are refined as needed to preserve perfect recall. The resulting game $\Gamma^{\prime}$ has exactly two players (player 1 and the merged opponent), while preserving the game tree, player 1’s information partitions, and player 1’s payoffs.
A key observation is that strict (or weak) dominance for player 1 is preserved under this transformation: an action is dominated in the original $n$-player game if and only if it is dominated in the merged two-player game. Therefore, once the game is transformed into $\Gamma^{\prime}$, we may directly apply the polynomial-time dominance tests developed above for the two-player case. In this way, both single-step dominance testing and iterated removal of dominated actions can be performed in polynomial time even in general $n$-player perfect-recall extensive-form games.
Let $\Gamma$ be a finite $n$-player extensive-form game of imperfect information with perfect recall, with players $N=\{1,2,\dots,n\}$. Fix player 1 and an information set $I_{1}$ of player 1. We construct a two-player game $\Gamma^{\prime}$ with players $1$ and $P$ (the “merged” opponent) as follows.
  • The underlying game tree (histories, chance nodes, terminal nodes, and payoffs) of $\Gamma^{\prime}$ is identical to that of $\Gamma$. In particular, the payoff function $u_{1}$ for player 1 is unchanged.
  • Every decision node controlled by player 1 in $\Gamma$ is still controlled by player 1 in $\Gamma^{\prime}$. Every decision node controlled by any player $j\in\{2,\dots,n\}$ in $\Gamma$ is controlled by player $P$ in $\Gamma^{\prime}$.
  • The information sets of player 1 in $\Gamma^{\prime}$ are exactly the same as in $\Gamma$.
  • The information sets of player $P$ are obtained from the information sets of players $2,\dots,n$ in $\Gamma$ by refinement to ensure perfect recall for $P$: for each original information set $J$ of some player $j\in\{2,\dots,n\}$, we partition $J$ into the equivalence classes induced by the relation $h\sim h^{\prime}$ if and only if the sequence of information sets and actions at which any of the original opponents $\{2,\dots,n\}$ moved along the path to $h$ is identical to that along the path to $h^{\prime}$.
By construction, $\Gamma^{\prime}$ is a two-player extensive-form game with perfect recall (for players $1$ and $P$), and its size is linear in the size of $\Gamma$.
Lemma 1.
Let $\Gamma$ be an $n$-player extensive-form game of imperfect information with perfect recall, and let $\Gamma^{\prime}$ be the two-player game obtained by merging players $2,\dots,n$ into a single opponent $P$ as above. Fix an information set $I_{1}$ of player 1 and an action $a\in A(I_{1})$. Then $a$ is strictly dominated for player 1 at $I_{1}$ in $\Gamma$ (according to Definition 1) if and only if it is strictly dominated for player 1 at $I_{1}$ in $\Gamma^{\prime}$.
Proof.
Let $\Sigma_{j}$ denote the set of behavior strategies of player $j$ in $\Gamma$, and let
$\Sigma_{-1}\;=\;\Sigma_{2}\times\cdots\times\Sigma_{n}$
denote the set of behavior-strategy profiles of all players other than 1. For an information set $I_{1}$ of player 1, let $\Sigma_{-1}^{I_{1}}\subseteq\Sigma_{-1}$ denote the subset of profiles that “always play to reach $I_{1}$ when possible,” as in Definition 1. Similarly, let $\Sigma_{P}$ denote the set of behavior strategies of player $P$ in $\Gamma^{\prime}$, and let $\Sigma_{P}^{I_{1}}\subseteq\Sigma_{P}$ be the subset that always plays to reach $I_{1}$ when possible in $\Gamma^{\prime}$.
Because behavior strategies assign an independent probability distribution to each information set, and because the information sets of $P$ in $\Gamma^{\prime}$ are obtained by refining and relabeling the information sets of players $2,\dots,n$ in $\Gamma$, there is a natural bijection
$\Phi:\Sigma_{-1}\;\longrightarrow\;\Sigma_{P}$
that maps a profile $(\sigma_{2},\dots,\sigma_{n})$ to a single behavior strategy $\sigma_{P}$ which, at each information set of some original player $j\geq 2$, plays according to $\sigma_{j}$. The local “always reach $I_{1}$” constraints are preserved under this mapping, so
$\Phi\big(\Sigma_{-1}^{I_{1}}\big)=\Sigma_{P}^{I_{1}}.$
For any profile $(\sigma_{1},\sigma_{-1})$ in $\Gamma$, and the corresponding $(\sigma_{1},\sigma_{P})$ in $\Gamma^{\prime}$ with $\sigma_{P}=\Phi(\sigma_{-1})$, the induced distribution over terminal histories is the same in $\Gamma$ and $\Gamma^{\prime}$, since the game tree and player 1’s actions are unchanged and the opponents’ local behavior at each information set is identical. Therefore, for all such profiles we have
$u_{1}(\sigma_{1},\sigma_{-1})\;=\;u_{1}(\sigma_{1},\sigma_{P}),$
where $u_{1}$ is player 1’s payoff.
Now fix an action $a\in A(I_{1})$. By Definition 1, $a$ is strictly dominated for player 1 at $I_{1}$ in $\Gamma$ if and only if there exists a behavior strategy $\sigma_{1}^{-a}$ for player 1 that always reaches $I_{1}$ and never plays $a$ at $I_{1}$ such that, for every behavior strategy $\sigma_{1}^{a}$ that always reaches $I_{1}$ and plays $a$ with probability 1 at $I_{1}$, and for every $\sigma_{-1}\in\Sigma_{-1}^{I_{1}}$,
$u_{1}(\sigma_{1}^{-a},\sigma_{-1})\;>\;u_{1}(\sigma_{1}^{a},\sigma_{-1}).$
Using the bijection $\Phi$ and equality of payoffs, this is equivalent to: there exists $\sigma_{1}^{-a}$ such that for every $\sigma_{1}^{a}$ and for every $\sigma_{P}\in\Sigma_{P}^{I_{1}}$,
$u_{1}(\sigma_{1}^{-a},\sigma_{P})\;>\;u_{1}(\sigma_{1}^{a},\sigma_{P}),$
which is precisely the strict-dominance condition for action $a$ at $I_{1}$ in the merged two-player game $\Gamma^{\prime}$ (again using Definition 1 with the opponents collapsed into $P$).
Thus $a$ is strictly dominated in $\Gamma$ if and only if it is strictly dominated in $\Gamma^{\prime}$. ∎
Theorem 4.
There exists a polynomial-time algorithm that determines whether a given action is strictly dominated in an $n$-player perfect-recall extensive-form game of imperfect information.
Proof.
Let $\Gamma$ be an $n$-player extensive-form game with perfect recall, let $I_{1}$ be an information set of player 1, and let $a\in A(I_{1})$ be the action to be tested.
First, construct the merged two-player game $\Gamma^{\prime}$ with players $1$ and $P$ as described above. This transformation does not duplicate any nodes of the game tree; it only relabels decision nodes and refines some information sets of the opponents. Hence the size of $\Gamma^{\prime}$ (number of nodes, information sets, and actions) is linear in the size of $\Gamma$, and the construction can be performed in polynomial time.
By Lemma 1, action $a$ is strictly dominated for player 1 at $I_{1}$ in $\Gamma$ (according to Definition 1) if and only if it is strictly dominated for player 1 at $I_{1}$ in $\Gamma^{\prime}$.
The game $\Gamma^{\prime}$ is a two-player extensive-form game with perfect recall. By the two-player result stated earlier, strict dominance of a given action in a two-player perfect-recall extensive-form game of imperfect information can be decided in time polynomial in the size of the game, using the sequence-form linear programs derived in the previous section.
Therefore, we can decide whether $a$ is strictly dominated in $\Gamma$ by:
  1. 1.
    constructing the merged game $\Gamma^{\prime}$ in polynomial time, and
  2. 2.
    running the two-player strict-dominance algorithm on $\Gamma^{\prime}$ in polynomial time.
The total running time is polynomial in the size of $\Gamma$, which proves the theorem. ∎
Theorem 5.
There exists a polynomial-time algorithm that determines whether a given action is weakly dominated for player 1 in an $n$-player perfect-recall extensive-form game of imperfect information.
Proof.
The weak-dominance test for two-player perfect-recall games is polynomial-time (Theorem 2). Weak dominance is preserved under the merging transformation exactly as in Lemma 1: the definition only modifies the payoff inequality from strict to weak, but the induced distributions and opponent strategy sets remain in bijection under the merge. Thus an action is weakly dominated in the original $n$-player game if and only if it is weakly dominated in the merged two-player game. Applying the two-player weak-dominance test to the merged game gives a polynomial-time algorithm for the $n$-player case. ∎
Theorem 6.
Iterated removal of strictly and weakly dominated actions can be performed in polynomial time in an $n$-player perfect-recall extensive-form game of imperfect information.
Proof.
At each iteration of the iterated-removal process in the $n$-player game $\Gamma$, for every information set $I_{1}$ of player 1 and every action $a\in A(I_{1})$, we test whether $a$ is strictly or weakly dominated according to Definitions 1 and 2. To perform this test, we construct the merged two-player game $\Gamma^{\prime}$ and apply the corresponding strict or weak-dominance algorithm for two-player games.
By Theorems 4 and 5, each dominance test in $\Gamma^{\prime}$ runs in time polynomial in the size of $\Gamma$. The number of actions in $\Gamma$ is polynomial in the size of the game tree, so in each iteration we perform only polynomially many dominance tests, and each such iteration therefore runs in polynomial time. Furthermore, each action can be removed at most once, so the total number of iterations is at most the total number of actions, and hence polynomial in the size of $\Gamma$.
Finally, Lemma 1 ensures that at every iteration the dominance status of each action is identical in the original $n$-player game $\Gamma$ and in the merged two-player game $\Gamma^{\prime}$. Therefore an action is removed in $\Gamma$ at a given iteration if and only if the corresponding action is removed in $\Gamma^{\prime}$. Combining these observations, we obtain that the entire iterated-removal process runs in polynomial time and yields the same reduced game whether it is viewed in $\Gamma$ or in $\Gamma^{\prime}$. ∎
6 Experiments
Now that we have defined dominated actions and showed that they can be computed efficiently, we would like to investigate whether they can be a useful concept in practice. Poker has been widely studied as a test domain for imperfect-information games. The most popular variant regularly played by humans is No-Limit Texas Hold’em (NLHE). Two-player NLHE works as follows. Initially two players each have a stack of chips. One player, called the small blind, initially puts $k$ worth of chips in the middle, while the other player, called the big blind, puts in $2k$. The chips in the middle are known as the pot, and will go to the winner of the hand. Next, there is an initial round of betting. The player whose turn it is to act can choose from three available options:
  • Fold: Give up on the hand, surrendering the pot to the opponent.
  • Call: Put in the minimum number of chips needed to match the number of chips put into the pot by the opponent. For example, if the opponent has put in \$1000 and we have put in \$400, a call would require putting in \$600 more. A call of zero chips is also known as a check.
  • Bet: Put in additional chips beyond what is needed to call. A bet can be of any size from 1 chip up to the number of chips a player has left in their stack, provided it exceeds some minimum value and is a multiple of the smallest chip denomination. A bet of all of one’s remaining chips is called an all-in bet (aka a shove).
The initial round of betting ends if a player has folded, if there has been a bet and a call, or if both players have checked. If the round ends without a player folding, then three public cards are revealed face-up on the table (called the flop) and a second round of betting takes place. Then one more public card is dealt (called the turn) and a third round of betting, followed by a fifth public card (called the river) and a final round of betting. If a player ever folds, the other player wins all the chips in the pot. If the final betting round is completed without a player folding (or if a player is all-in at an earlier round), then both players reveal their private cards, and the player with the best five-card hand (out of their two private cards and the five public cards) wins the pot (it is divided equally if there is a tie).
In some situations the blinds are very large relative to stack sizes. This can happen frequently at later stages in poker tournaments, where the blinds increase after a certain time duration. A common rule is that when stack sizes are below around 8 big blinds a shove-or-fold strategy should be employed where each player only goes all-in or folds [10]. Study of optimal shove-or-fold strategies has been considered for 2-player [11] and 3-player [4, 5] poker tournament endgames. The poker site Americas Cardroom [^2] https://www.americascardroom.eu/ has specific “All-in or Fold” tables with up to 4 players where players are only allowed to play shove-or-fold strategies. The initial stack sizes at these tables are either 8 or 10 times the big blind; the highest stake available has blinds of \$100 and \$200 with initial stacks of \$2000.
In the two-player NLHE shove-or-fold game, each player has 169 strategically distinct hands with which they can choose to shove or fold (13 pocket pairs and $\frac{13\cdot 12}{2}=78$ combinations of each non-paired offsuit and suited hand). Let player 1 denote the small blind and player 2 denote the big blind. We assume that the blinds are small blind $k=100$ and big blind $2k=200$, and initial stacks are 1600 (8 times the big blind). We first remove all strictly dominated actions for player 1, followed by removing strictly dominated actions for player 2 (note that removing weakly dominated actions does not provide additional benefit in this game). It turns out that 85 actions for player 1 are removed and 99 actions for player 2 are removed. Thus, the initial game with 169 hands per player can be reduced to a game where player 1 must make a decision with 84 hands and player 2 must make a decision with 70 hands; the number of decision points has been reduced by over 50%. It turns out that performing an additional round of removing dominated actions does not remove any further actions. [^3] We note that no actions are strongly dominated for either player in this game (i.e., Candidate Definition 1). For example, consider the action of player 1 folding with pocket aces (the best starting hand). The payoff of this action is -\$100. If player 1 goes all-in, there are leaf nodes where player 1 will receive a payoff of -\$1600.
Next we consider the setting where the stacks are 5 times the big blind (i.e., 1000). In this game five rounds of iterated removal of dominated actions are needed. The first round removes 108 dominated actions for player 1 and 129 for player 2; the second round removes 20 for player 1 and 16 for player 2; the third round removes 8 for player 1 and 6 for player 2; the fourth round removes 7 for player 1 and 2 for player 2; the fifth round removes 1 for player 1 and 0 for player 2. In the final reduced game player 1 must make a decision with only 25 hands while player 2 must make a decision with only 16 hands. Tables 1 and 2 show the non-dominated actions for each player with parentheses indicating the iteration at which the alternative action was removed. ‘S’ indicated shove, ‘F’ indicates fold, and ‘?’ indicates that neither action was dominated. If stacks are 4 times the big blind the game is solved completely after 4 rounds of removing dominated actions, and for stacks of 3 big blinds the game is solved completely after 2 rounds. These results demonstrate that iteratively removing dominated actions can significantly reduce the size of realistic games. While for simplicity we considered two-player zero-sum games, for which the full game can be solved directly by a linear program, the computational benefit for games with more than two players could be much more significant.
A K Q J T 9 8 7 6 5 4 3 2
A S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1)
K S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1)
Q S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1)
J S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1)
T S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(3) ?
9 S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(3) ? ? ?
8 S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(2) S(4) ? ? ?
7 S(1) S(1) S(1) S(1) S(1) S(3) S(5) S(1) S(2) S(3) ? ? F(3)
6 S(1) S(1) S(1) S(1) ? ? ? ? S(1) S(3) ? ? F(4)
5 S(1) S(1) S(1) S(1) ? ? ? ? ? S(1) S(4) ? ?
4 S(1) S(1) S(1) S(2) F(4) F(2) F(2) F(3) F(4) ? S(1) ? F(4)
3 S(1) S(1) S(1) ? F(2) F(2) F(2) F(2) F(2) F(2) F(2) S(1) F(3)
2 S(1) S(1) S(1) F(4) F(2) F(2) F(2) F(2) F(2) F(2) F(2) F(2) S(1)
Table 1: Dominated actions for player 1 in 2-player NLHE allin-fold with 5 big blind stacks. Suited hands are in the upper right and unsuited hands are in the lower left.
A K Q J T 9 8 7 6 5 4 3 2
A S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1)
K S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1)
Q S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1)
J S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(3) S(3) ?
T S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(2) ? ? ? ?
9 S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(1) S(3) ? F(4) F(3) F(2)
8 S(1) S(1) S(1) S(1) S(1) S(3) S(1) S(1) S(1) ? F(2) F(2) F(1)
7 S(1) S(1) S(1) S(2) ? ? ? S(1) S(1) S(1) F(2) F(1) F(1)
6 S(1) S(1) S(1) ? ? ? F(3) F(2) S(1) S(1) F(2) F(1) F(1)
5 S(1) S(1) S(1) ? F(4) F(2) F(2) F(2) F(1) S(1) S(1) F(1) F(1)
4 S(1) S(1) S(1) ? F(2) F(2) F(1) F(1) F(1) F(1) S(1) F(1) F(1)
3 S(1) S(1) S(1) ? F(2) F(1) F(1) F(1) F(1) F(1) F(1) S(1) F(1)
2 S(1) S(1) S(1) F(2) F(2) F(1) F(1) F(1) F(1) F(1) F(1) F(1) S(1)
Table 2: Dominated actions for player 2 in 2-player NLHE allin-fold with 5 big blind stacks. Suited hands are in the upper right and unsuited hands are in the lower left.
7 Conclusion
Dominance is a fundamental concept in game theory. It is well-understood in normal-form games, but its impact in imperfect-information games has so far been unexplored. We consider several plausible definitions of dominated actions which we demonstrate to be problematic; however, one of them which we denote as strong domination can be useful as an efficient preprocessing step. We present a new definition that addresses these limitations. We show that both strictly and weakly dominated actions can be identified in polynomial time, and that iterative removal can be performed in polynomial time in $n$-player games. Our algorithms identify actions that are dominated by any mixed strategy, not necessarily a pure action. We demonstrate that removing dominated actions can play a significant role in reducing the size of realistic imperfect-information games. This can serve as an efficient preprocessing step before computation of a Nash equilibrium. Subsequent recent work has shown that removal of dominated actions enabled an algorithm to compute a Nash equilibrium in a three-player imperfect-information game in under 3 seconds while the algorithm was unable to solve the full game in 24 hours [3]. In practice our algorithms could be sped up by several heuristics such as selection of the order of traversal of information sets and players. Recent work has shown that some games contain many “mistake” actions that are played with probability zero in all Nash equilibria but are not dominated [2]. Thus, there is potentially more future ground to explore on efficient game reduction by elimination of poor actions.

### v7

@@
 In this paper we define and study the concept of dominated actions in imperfect-information games.
-Our main result is a polynomial-time algorithm for determining whether an action is dominated (strictly or weakly) by any mixed strategy in $n$-player games, which can be extended to an algorithm for iteratively removing dominated actions.
+Our main result is a polynomial-time algorithm for determining whether an action is dominated (strictly or weakly) by any mixed strategy in two-player perfect-recall games with publicly observable actions, which can be extended to iteratively remove dominated actions.
 This allows us to efficiently reduce the size of the game tree as a preprocessing step for Nash equilibrium computation.
@@
 (Note that the requirement that this holds for all opposing pure strategy profiles is enough to ensure that it also holds for all mixed strategy profiles of the opponents as well.) Clearly it would be irrational for an agent to play a strictly dominated strategy, as another strategy will do strictly better regardless of the beliefs of what the opponents would play.
-Strategy $\sigma^{\prime}_{i}$ weakly dominates $\sigma$ if the inequality holds weakly (though strictly for at least one pure strategy profile): formally, $u_{i}(\sigma^{\prime},s_{-i})\geq u_{i}(\sigma_{i},s_{-i})$ for all $s_{-i}\in S_{-i}$ where the inequality is strict for at least one $s_{-i}$.
+Strategy $\sigma^{\prime}_{i}$ weakly dominates $\sigma_{i}$ if the inequality holds weakly (though strictly for at least one pure strategy profile): formally, $u_{i}(\sigma^{\prime}_{i},s_{-i})\geq u_{i}(\sigma_{i},s_{-i})$ for all $s_{-i}\in S_{-i}$ where the inequality is strict for at least one $s_{-i}$.
 (The condition that at least one inequality is strict is simply to rule out saying that a strategy is dominated by an identical strategy).
@@
 This algorithm simply iterates over each pure strategy $s_{i}$ for player $i$ and tests whether it performs strictly better than $\sigma_{i}$ against each opposing pure strategy profile $s_{-i}$.
-This procedure has complexity $O(|A|)$, where $A=\times_{i}A_{i}$ is the set of joint pure strategy profiles, and so takes time linear in the size of the game.
+This procedure has complexity $O(|S|)$, where $S=\times_{i}S_{i}$ is the set of joint pure strategy profiles, and so takes time linear in the size of the game.
 The procedure can also be easily adapted to produce an algorithm for determining whether a mixed strategy profile is weakly dominated by any pure strategy in linear time.
@@
 These characterizations, however, do not extend to extensive-form games.
-In imperfect-information settings, the dominance of an action depends on which opponent strategies reach the corresponding information set, a feature with no analogue in the normal-form framework.
+In imperfect-information settings, the dominance of an action depends on which opponent strategies assign positive probability to histories reaching the corresponding information set, a feature with no analogue in the normal-form framework.
 Consequently, an action in an extensive-form game may be locally dominated even when every pure or mixed strategy containing that action is undominated in the induced normal form.
@@
 If no player forgets information that they previously knew, we say that the game has perfect recall.
+A game has publicly observable actions if, for every player, no information set contains nodes reached by different actions of another player.
 A (behavioral) strategy for player $i,$ $\sigma_{i}\in\Sigma_{i},$ is a function that assigns a probability distribution over all actions at each information set belonging to $i$.
 In theory, every extensive-form game can be converted to an equivalent normal-form game; however, there is an exponential blowup in the size of the game representation, and therefore such a conversion is undesirable.
-Instead, new algorithms have been developed that operate on the extensive form representation directly.
+Instead, algorithms have been developed that operate on the extensive form representation directly.
 It turns out that the complexity of computing equilibria in extensive-form games is similar to that of normal-form games; a Nash equilibrium can be computed in polynomial time in two-player zero-sum games (with perfect recall) [8], while the problem is hard for two-player general-sum and multiplayer games.
@@
 The method uses several matrices defined as follows.
-For player 1, the matrix $\mathbf{E}$ is defined where each row corresponds to an information set (including an initial row for the “empty” information set), and each column corresponds to an action sequence (including an initial row for the “empty” action sequence).
+For player 1, the matrix $\mathbf{E}$ is defined where each row corresponds to an information set (including an initial row for the “empty” information set), and each column corresponds to an action sequence (including an initial column for the “empty” action sequence).
 In the first row of $\mathbf{E}$ the first element is 1 and all other elements are 0; subsequent rows have -1 for the entries corresponding to the action sequence leading to the root of the information set, and 1 for all actions that can be taken at the information set (and 0 otherwise).
-Thus $\mathbf{E}$ has dimension $c_{1}\times d_{1}$, where $c_{i}$ is the number of information sets for player $i$ and $d_{i}$ is the number of action sequences for player $i$.
+Thus $\mathbf{E}$ has dimension $(c_{1}+1)\times(d_{1}+1)$, where $c_{i}$ is the number of information sets for player $i$ and $d_{i}$ is the number of action sequences for player $i$.
 Matrix $\mathbf{F}$ is defined analogously for player 2.
-The vector $\mathbf{e}$ is defined to be a column vector of length $c_{1}$ with 1 in the first position and 0 in other entries, and vector $\mathbf{f}$ is defined with length $c_{2}$ analogously.
-The matrix $\mathbf{A}$ is defined with dimension $d_{1}\times d_{2}$ where entry $A_{ij}$ gives the payoff for player 1 when player 1 plays action sequence $i$ and player 2 plays action sequence $j$ multiplied by the probabilities of chance moves along the path of play.
+The vector $\mathbf{e}$ is defined to be a column vector of length $c_{1}+1$ with 1 in the first position and 0 in other entries, and vector $\mathbf{f}$ is defined with length $c_{2}+1$ analogously.
+The matrix $\mathbf{A}$ is defined with dimension $(d_{1}+1)\times(d_{2}+1)$ where entry $A_{ij}$ gives the payoff for player 1 when player 1 plays action sequence $i$ and player 2 plays action sequence $j$ multiplied by the probabilities of chance moves along the path of play.
 The matrix $\mathbf{B}$ of player 2’s expected payoffs is defined analogously.
@@
 Given these matrices we can solve one of two linear programming problems to compute a Nash equilibrium in zero-sum extensive-form games [8].
-In the first formulation the primal variables $\mathbf{x}$ correspond to player 1’s mixed strategy while the dual variables correspond to player 2’s strategy.
-In the second formulation, which is the dual problem of the first formulation, the primal decision variables $\mathbf{y}$ correspond to player 2’s strategy while the dual variables correspond to player 1’s strategy.
+In the first formulation the primal variables $\mathbf{x}$ correspond to player 1’s strategy while the dual variables $\mathbf{q}$ correspond to player 2’s strategy.
+In the second formulation, which is the dual problem of the first formulation, the primal decision variables $\mathbf{y}$ correspond to player 2’s strategy while the dual variables $\mathbf{p}$ correspond to player 1’s strategy.
 […]
@@
 Since player 2 does not take any actions, player 1 always achieves a strictly higher expected utility by playing action 1.
-However, action 2 is not strictly dominated according to Candidate Definition 1 because in one leaf node succeeding action 1 player 1 obtains payoff -100 which is lower than the payoff of -50 at leaf nodes following action 2.
+However, action 2 is not strictly dominated according to Candidate Definition 1 because in one leaf node succeeding action 1 player 1 obtains payoff 100-100 which is lower than the payoff of 50-50 at leaf nodes following action 2.
 Figure 1: Example two-player imperfect-information extensive-form game.
@@
 Suppose we want to apply Candidate Definition 3 to determine whether action 2 is strictly dominated for player 2.
-Then $\sigma^{-a_{i}}_{i}$ is the strategy for player 2 that plays action 1 with probability 1, and $\sigma^{a_{i}}_{i}$ must be the strategy for player 2 that plays action 2 with probability 1.
+Then $\sigma^{-a_{i}}_{i}$ is the strategy for player 2 that always plays action 1, and $\sigma^{a_{i}}_{i}$ must be the strategy for player 2 that plays action 2 with probability 1.
 Now suppose that player 1 plays action 2 with probability 1, which we denote as $\sigma_{-i}$.
@@
 Candidate Definition 3.
-Action $a_{i}$ for player $i$ is strictly dominated at information set $I_{i}$ if there exists a mixed strategy $\sigma^{-a_{i}}_{i}$ that plays action $a_{i}$ at $I_{i}$ with probability 0 such that for every mixed strategy $\sigma^{a_{i}}_{i}$ for player $i$ that plays action $a_{i}$ with probability 1 at $I_{i}$, $u_{i}\left(\sigma^{-a_{i}}_{i},\sigma_{-i}\right)>u_{i}\left(\sigma^{a_{i}}_{i},\sigma_{-i}\right)$ for all opposing strategies $\sigma_{-i}\in\Sigma_{-i}.$
+Action $a_{i}$ for player $i$ is strictly dominated at information set $I_{i}$ if there exists a behavioral strategy $\sigma^{-a_{i}}_{i}$ that plays action $a_{i}$ at $I_{i}$ with probability 0 such that for every behavioral strategy $\sigma^{a_{i}}_{i}$ for player $i$ that plays action $a_{i}$ with probability 1 at $I_{i}$, $u_{i}\left(\sigma^{-a_{i}}_{i},\sigma_{-i}\right)>u_{i}\left(\sigma^{a_{i}}_{i},\sigma_{-i}\right)$ for all opposing strategies $\sigma_{-i}\in\Sigma_{-i}.$
 Figure 3: Extensive-form game demonstrating problem with Candidate Definition 3.
@@
 Strictly-dominated actions are defined in Definition 1 and weakly-dominated actions are defined in Definition 2.
-In these definitions $\Sigma^{I_{i}}_{-i}$ denotes the set of mixed strategy profiles for the opponents that always take actions leading to information set $I_{i}$ when possible.
+In these definitions $\Sigma^{I_{i}}_{-i}$ denotes the set of behavioral strategy profiles for the opponents that select only actions that do not prevent reaching $I_{i}$ whenever such actions exist.
 Note that these definitions apply to games with any number of players.
-They also consider actions that are dominated by any mixed strategy (not necessarily just a pure action at $I_{i}$).
+They also consider actions that are dominated by any behavioral strategy (not necessarily just a pure action at $I_{i}$).
 It is easy to verify that these definitions address the issues that arose in the examples from Figure 1 and Figure 3.
@@
 Definition 1.
-Action $a_{i}$ for player $i$ is strictly dominated at information set $I_{i}$ if there exists a mixed strategy $\sigma^{-a_{i}}_{i}$ that always plays to get to $I_{i}$ and plays action $a_{i}$ at $I_{i}$ with probability 0 such that for every mixed strategy $\sigma^{a_{i}}_{i}$ for player $i$ that always plays to get to $I_{i}$ and plays action $a_{i}$ with probability 1 at $I_{i}$, $u_{i}\left(\sigma^{-a_{i}}_{i},\sigma_{-i}\right)>u_{i}\left(\sigma^{a_{i}}_{i},\sigma_{-i}\right)$ for all opposing strategies $\sigma_{-i}\in\Sigma^{I_{i}}_{-i}.$
+Action $a_{i}$ for player $i$ is strictly dominated at information set $I_{i}$ if there exists a behavioral strategy $\sigma^{-a_{i}}_{i}$ that always plays to get to $I_{i}$ and plays action $a_{i}$ at $I_{i}$ with probability 0 such that for every behavioral strategy $\sigma^{a_{i}}_{i}$ for player $i$ that always plays to get to $I_{i}$ and plays action $a_{i}$ with probability 1 at $I_{i}$, $u_{i}\left(\sigma^{-a_{i}}_{i},\sigma_{-i}\right)>u_{i}\left(\sigma^{a_{i}}_{i},\sigma_{-i}\right)$ for all opposing strategies $\sigma_{-i}\in\Sigma^{I_{i}}_{-i}.$
 Definition 2.
-Action $a_{i}$ for player $i$ is weakly dominated at information set $I_{i}$ if there exists a mixed strategy $\sigma^{-a_{i}}_{i}$ that always plays to get to $I_{i}$ and plays action $a_{i}$ at $I_{i}$ with probability 0 such that for every mixed strategy $\sigma^{a_{i}}_{i}$ for player $i$ that always plays to get to $I_{i}$ and plays action $a_{i}$ with probability 1 at $I_{i}$, $u_{i}\left(\sigma^{-a_{i}}_{i},\sigma_{-i}\right)\geq u_{i}\left(\sigma^{a_{i}}_{i},\sigma_{-i}\right)$ for all opposing strategies $\sigma_{-i}\in\Sigma^{I_{i}}_{-i}$, where the inequality is strict for at least one $\sigma_{-i}.$
+Action $a_{i}$ for player $i$ is weakly dominated at information set $I_{i}$ if there exists a behavioral strategy $\sigma^{-a_{i}}_{i}$ that always plays to get to $I_{i}$ and plays action $a_{i}$ at $I_{i}$ with probability 0 such that for every behavioral strategy $\sigma^{a_{i}}_{i}$ for player $i$ that always plays to get to $I_{i}$ and plays action $a_{i}$ with probability 1 at $I_{i}$, $u_{i}\left(\sigma^{-a_{i}}_{i},\sigma_{-i}\right)\geq u_{i}\left(\sigma^{a_{i}}_{i},\sigma_{-i}\right)$ for all opposing strategies $\sigma_{-i}\in\Sigma^{I_{i}}_{-i}$, where the inequality is strict for at least one $\sigma_{-i}.$
 The example games considered were created using the open-source software package Gambit [13].
@@
 4 Algorithm for identifying dominated actions in two-player games
-Suppose we want to determine whether an action $c$ for player 1 is dominated at information set $I$ in a two-player perfect-recall extensive-form game $G$.
-Using the sequence form representation, suppose that action $c$ is the final action in the sequence with index $i$ for player 1.
-Let $S_{I}$ denote the set of indices corresponding to action sequences leading to $I$.
-We would like to solve the following problem.
+Suppose we want to determine whether an action $c$ for player 1 is dominated at information set $I$ in a two-player perfect-recall extensive-form game with publicly observable actions $G$.
+Using the sequence form representation, suppose that action $c$ is the final action in the sequence with index $i_{c}$ for player 1, and suppose that $i_{p}$ is the index of the parent sequence to sequence $i_{c}.$ Let $S_{I}$ denote the set of opponent sequences that appear on at least one path from the root to a node in information set $I.$ We restrict the opponent strategy vector so that only sequences in $S_{I}$ can receive positive probability.
+Each sequence corresponds to a component of the realization vector; we refer to the component index associated with sequences $s$ as its index.
+To consider whether action $c$ is dominated, we consider the following sequence-form optimization problem.
 […]
@@
 Proof.
-Let $f_{1}$ be the optimal objective value in Problem 1 and $f_{2}$ be the optimal objective value in Problem 2.
-Suppose that the optimal variables in Problem 1 are $\mathbf{x}^{1}_{1},\mathbf{x}^{1}_{2},\mathbf{y}^{1}.$ Now set $\mathbf{x}^{2}_{1}=\mathbf{x}^{1}_{1},$ $\mathbf{x}^{2}_{2}=\mathbf{x}^{1}_{2},$ $\mathbf{y}^{2}_{1}=\mathbf{y}^{1},$ $\mathbf{y}^{2}_{2}=\mathbf{y}^{1}.$ Then $(\mathbf{x}^{2}_{1},\mathbf{x}^{2}_{2},\mathbf{y}^{2}_{1},\mathbf{y}^{2}_{2})$ gives a feasible solution to Problem 2 with objective value equal to $f_{1}.$ So $f_{2}\geq f_{1}$.
-Now suppose that the optimal variables in Problem 2 are $\mathbf{x}^{2}_{1},\mathbf{x}^{2}_{2},\mathbf{y}^{2}_{1},\mathbf{y}^{2}_{2}.$ Now set $\mathbf{x}^{1}_{1}=\mathbf{x}^{2}_{1},$ $\mathbf{x}^{1}_{2}=\mathbf{x}^{2}_{2},$ and set $\mathbf{y}^{1}$ equal to the strategy that follows $\mathbf{y}^{2}_{1}$ at states following action $c$ for player 1, and follows $\mathbf{y}^{2}_{2}$ otherwise.
-Then $(\mathbf{x}^{1}_{1})^{T}\mathbf{A}\mathbf{y}^{1}=(\mathbf{x}^{1}_{1})^{T}\mathbf{A}\mathbf{y}^{2}_{1},$ since both players only take actions to get to information set $I$, $\mathbf{x}^{1}_{1}$ takes action $c$ at $I$, and the strategies $\mathbf{y}^{1}$ and $\mathbf{y}^{2}_{1}$ are identical after player 1 takes action $c$ at $I.$ Similarly, $(\mathbf{x}^{1}_{2})^{T}\mathbf{A}\mathbf{y}^{1}=(\mathbf{x}^{1}_{2})^{T}\mathbf{A}\mathbf{y}^{2}_{2},$ since both players only take actions to get to $I$, $\mathbf{x}^{1}_{1}$ does not take action $c$ at $I$, and the strategies $\mathbf{y}^{1}$ and $\mathbf{y}^{2}_{2}$ are identical after player 1 does not take action $c$ at $I.$ So $f_{1}\geq f_{2}$.
-So we conclude that $f_{1}=f_{2}.$
+For a fixed feasible $\mathbf{x}_{2}$, let
+$\phi_{1}(\mathbf{x}_{2})=\min_{\mathbf{x}_{1},\mathbf{y}}\left(\mathbf{x}_{2}^{T}A\mathbf{y}-\mathbf{x}_{1}^{T}A\mathbf{y}\right)$
+denote the inner objective value of Problem 1, and let
+$\phi_{2}(\mathbf{x}_{2})=\min_{\mathbf{x}_{1},\mathbf{y}_{1},\mathbf{y}_{2}}\left(\mathbf{x}_{2}^{T}A\mathbf{y}_{2}-\mathbf{x}_{1}^{T}A\mathbf{y}_{1}\right)$
+denote the inner objective value of Problem 2.
+First observe that $\phi_{2}(\mathbf{x}_{2})\leq\phi_{1}(\mathbf{x}_{2})$ for every feasible $\mathbf{x}_{2}$, since any feasible pair $(\mathbf{x}_{1},\mathbf{y})$ for Problem 1 yields a feasible triple $(\mathbf{x}_{1},\mathbf{y}_{1},\mathbf{y}_{2})$ for Problem 2 by setting $\mathbf{y}_{1}=\mathbf{y}_{2}=\mathbf{y}$.
+Now fix a feasible $\mathbf{x}_{2}$ and let $(\mathbf{x}_{1},\mathbf{y}_{1},\mathbf{y}_{2})$ be any feasible solution of the inner minimization in Problem 2.
+Define a strategy $\mathbf{y}$ that follows $\mathbf{y}_{1}$ at all information sets reachable after player 1 takes action $c$ at information set $I$, and follows $\mathbf{y}_{2}$ otherwise.
+Here we use the assumption that actions are publicly observable.
+Because player 1’s action at $I$ is publicly observed, no information set of player 2 can contain nodes reached both after action $c$ and after a different action at $I$.
+Therefore, the prescriptions of $\mathbf{y}_{1}$ and $\mathbf{y}_{2}$ are never assigned inconsistently at the same information set, so this stitching defines a valid behavioral strategy $\mathbf{y}$ for Problem 1.
+Moreover,
+$\mathbf{x}_{1}^{T}A\mathbf{y}=\mathbf{x}_{1}^{T}A\mathbf{y}_{1},\qquad\mathbf{x}_{2}^{T}A\mathbf{y}=\mathbf{x}_{2}^{T}A\mathbf{y}_{2},$
+because $\mathbf{x}_{1}$ takes action $c$ at $I$ while $\mathbf{x}_{2}$ does not. Hence
+$\mathbf{x}_{2}^{T}A\mathbf{y}-\mathbf{x}_{1}^{T}A\mathbf{y}=\mathbf{x}_{2}^{T}A\mathbf{y}_{2}-\mathbf{x}_{1}^{T}A\mathbf{y}_{1}.$
+Since this holds for every feasible $(\mathbf{x}_{1},\mathbf{y}_{1},\mathbf{y}_{2})$, we obtain $\phi_{1}(\mathbf{x}_{2})\leq\phi_{2}(\mathbf{x}_{2})$.
+Thus $\phi_{1}(\mathbf{x}_{2})=\phi_{2}(\mathbf{x}_{2})$ for every feasible $\mathbf{x}_{2}$.
+Maximizing over $\mathbf{x}_{2}$ in the two problems yields $v_{1}=v_{2}$.
+∎
 Proposition 1 allows us to divide Problem 2 into the following two subproblems.
-If $f$ is the optimal objective value of Problem 2, $f_{1}$ is the optimal objective of Problem 3, and $f_{2}$ is the optimal objective of Problem 4, then we have $f=f_{1}-f_{2}.$
+If $v_{2}$ is the optimal objective value of Problem 2, $v_{3}$ is the optimal objective of Problem 3, and $v_{4}$ is the optimal objective of Problem 4, then we have $v_{2}=v_{3}-v_{4}.$
 […]
@@
 The Lagrangian is
-$L(\mathbf{y},\boldsymbol{\lambda},\boldsymbol{\gamma},\mathbf{r})=\mathbf{x}^{T}_{2}\mathbf{A}\mathbf{y}-(\mathbf{f}^{T}-\mathbf{y}^{T}\mathbf{F}^{T})\boldsymbol{\lambda}-\sum_{k\in S_{I}}\boldsymbol{\gamma}_{k}(\mathbf{y}_{k}-1)-\mathbf{r}^{T}\mathbf{y}$
-$\frac{\partial L}{\partial\mathbf{y}}=\mathbf{x}^{T}_{2}\mathbf{A}+\boldsymbol{\lambda}^{T}\mathbf{F}-\sum_{k\in S_{I}}\boldsymbol{\gamma}_{k}\mathbf{e_{k}}-\mathbf{r}^{T}$
+$L(\mathbf{y},\boldsymbol{\lambda},\boldsymbol{\gamma},\mathbf{r})=\mathbf{x}^{T}_{2}\mathbf{A}\mathbf{y}-(\mathbf{f}^{T}-\mathbf{y}^{T}\mathbf{F}^{T})\boldsymbol{\lambda}-\sum_{k\notin S_{I}}\boldsymbol{\gamma}_{k}\mathbf{y}_{k}-\mathbf{r}^{T}\mathbf{y}$
+$\frac{\partial L}{\partial\mathbf{y}}=\mathbf{A}^{T}\mathbf{x}_{2}+\mathbf{F}^{T}\boldsymbol{\lambda}-\sum_{k\notin S_{I}}\boldsymbol{\gamma}_{k}\mathbf{e_{k}}-\mathbf{r}$
 The dual problem is
@@
 Note that both players are aligned in trying to maximize the objective.
-Let us define a new problem $\overline{G}$ where player 1 selects the actions for both players.
-We denote player 1’s strategy in this modified game by $\overline{\mathbf{x}}_{1}.$ We modify $\mathbf{E}$ and resize $\mathbf{e}$ accordingly and denote them by $\mathbf{E^{\prime}}$ and $\mathbf{e^{\prime}}.$ The payoffs can now be represented as a vector $\mathbf{a}$.
-In the new representation let $i^{\prime}$ denote the index of the sequence with concluding action $c$, and let $S^{\prime}_{I}$ denote the set of indices corresponding to action sequences leading to $I$.
+Let us define a new problem $G^{\prime}$ where player 1 selects the actions for both players.
+We denote player 1’s strategy in this modified game by $\mathbf{x}^{\prime}_{1}.$ We modify $\mathbf{E}$ and resize $\mathbf{e}$ accordingly and denote them by $\mathbf{E^{\prime}}$ and $\mathbf{e^{\prime}}.$ The payoffs can now be represented as a vector $\mathbf{a}^{\prime}$.
+In the new representation let $i^{\prime}_{c}$ denote the index of the sequence with concluding action $c$, and let $i^{\prime}_{p}$ denote the index of the parent sequence.
 […]
-Let $u_{2}$ denote the optimal objective value for optimization problem 5, and $u_{1}$ denote the optimal objective value for problem 6.
-If $u_{2}>u_{1}$ then we conclude that action $c$ is strictly dominated.
-If $u_{2}<u_{1}$ then we conclude that action $c$ is not strictly or weakly dominated.
-If $u_{2}=u_{1}$, let $u_{3}$ denote the optimal objective value for problem 7, and let $u_{4}$ denote the optimal objective value for problem 8.
-If $u_{3}>u_{4}$ then we conclude that action $c$ is weakly dominated, and if $u_{3}=u_{4}$ we conclude that action $c$ is not strictly or weakly dominated (note that we cannot have $u_{3}<u_{4}$).
+Note that $G^{\prime}$ is not treated as a two-player game and no equilibrium computation is performed.
+Instead, player 1 controls all decision nodes and the problem becomes a single-agent optimization problem over the game tree.
+The sequence-form constraints $E^{\prime}$ simply enforce valid realization plans in the tree and do not rely on perfect recall.
+Therefore the formulation remains valid even though merging the players may introduce imperfect recall.
+The same interpretation applies to Problems 7 and 8.
+Let $v_{5}$ denote the optimal objective value for optimization problem 5, and $v_{6}$ denote the optimal objective value for problem 6.
+If $v_{5}>v_{6}$ then we conclude that action $c$ is strictly dominated.
+If $v_{5}<v_{6}$ then we conclude that action $c$ is not strictly or weakly dominated.
+If $v_{5}=v_{6}$, let $v_{7}$ denote the optimal objective value for problem 7, and let $v_{8}$ denote the optimal objective value for problem 8.
+If $v_{7}>v_{8}$ then we conclude that action $c$ is weakly dominated, and if $v_{7}=v_{8}$ we conclude that action $c$ is not strictly or weakly dominated (note that we cannot have $v_{7}<v_{8}$).
 […]
@@
 We can perform this procedure for every action at each information set of player 1, and analogously for player 2.
+The procedure is summarized in Algorithm 1.
 Since the number of actions is linear in the size of the game tree, the overall procedure involves solving a linear number of linear programs and therefore runs in polynomial time.
@@
 Note that the procedure applies to all two-player games and does not assume that they are zero sum.
-The procedure also removes actions that are dominated by any mixed strategy (which may play a probability distribution over actions at the same information set $I$), not just actions that are dominated by a pure action.
+The procedure also removes actions that are dominated by any behavioral strategy (which may play a probability distribution over actions at the same information set $I$), not just actions that are dominated by a pure action.
+Algorithm 1 Dominance test for action $c$ at information set II
+1: Construct the sequence-form matrices $(\mathbf{A},\mathbf{E},\mathbf{F})$ for game $G$
+2: Let $i_{c}$ denote the sequence index ending in action $c$, and let $i_{p}$ denote the parent sequence
+3: Compute $S_{I}$, the set of opponent sequences that reach information set II
+4: Solve LP (5) and obtain value $v_{5}$
+5: Solve LP (6) and obtain value $v_{6}$
+6:if $v_{5}>v_{6}$ then
+7:  return action $c$ is strictly dominated
+8:else if $v_{5}<v_{6}$ then
+9:  return action $c$ is not dominated
+10:end if
+11: Solve LP (7) and obtain value $v_{7}$
+12: Solve LP (8) and obtain value $v_{8}$
+13:if $v_{7}>v_{8}$ then
+14:  return action $c$ is weakly dominated
+15:else
+16:  return action $c$ is not dominated
+17:end if
 Theorem 1.
-There exists a polynomial-time algorithm that determines whether a given action is strictly dominated in a two-player perfect-recall extensive-form game of imperfect information.
+There exists a polynomial-time algorithm that determines whether a given action is strictly dominated in a two-player perfect-recall extensive-form game of imperfect information with publicly observable actions.
 Theorem 2.
-There exists a polynomial-time algorithm that determines whether a given action is weakly dominated in a two-player perfect-recall extensive-form game of imperfect information.
+There exists a polynomial-time algorithm that determines whether a given action is weakly dominated in a two-player perfect-recall extensive-form game of imperfect information with publicly observable actions.
 Theorem 3.
-Iterated removal of strictly and weakly dominated actions can be performed in polynomial time in a two-player perfect-recall extensive-form game of imperfect information.
-5 Extension to $n$ players
-Suppose we wish to determine whether an action of player 1 is strictly or weakly dominated in an $n$-player perfect-recall game $\Gamma$ with $n>2$.
-Although the linear-programming characterization developed above applies only to two-player games with perfect recall, we can reduce the general case to the two-player setting by merging all players other than player 1 into a single aggregated opponent.
-This merged player controls every decision node that originally belonged to players $2,\dots,n$, and its information sets are refined as needed to preserve perfect recall.
-The resulting game $\Gamma^{\prime}$ has exactly two players (player 1 and the merged opponent), while preserving the game tree, player 1’s information partitions, and player 1’s payoffs.
-A key observation is that strict (or weak) dominance for player 1 is preserved under this transformation: an action is dominated in the original $n$-player game if and only if it is dominated in the merged two-player game.
-Therefore, once the game is transformed into $\Gamma^{\prime}$, we may directly apply the polynomial-time dominance tests developed above for the two-player case.
-In this way, both single-step dominance testing and iterated removal of dominated actions can be performed in polynomial time even in general $n$-player perfect-recall extensive-form games.
-Let $\Gamma$ be a finite $n$-player extensive-form game of imperfect information with perfect recall, with players $N=\{1,2,\dots,n\}$.
-Fix player 1 and an information set $I_{1}$ of player 1.
-We construct a two-player game $\Gamma^{\prime}$ with players $1$ and $P$ (the “merged” opponent) as follows.
-  • The underlying game tree (histories, chance nodes, terminal nodes, and payoffs) of $\Gamma^{\prime}$ is identical to that of $\Gamma$.
-In particular, the payoff function $u_{1}$ for player 1 is unchanged.
-  • Every decision node controlled by player 1 in $\Gamma$ is still controlled by player 1 in $\Gamma^{\prime}$.
-Every decision node controlled by any player $j\in\{2,\dots,n\}$ in $\Gamma$ is controlled by player $P$ in $\Gamma^{\prime}$.
-  • The information sets of player 1 in $\Gamma^{\prime}$ are exactly the same as in $\Gamma$.
-  • The information sets of player $P$ are obtained from the information sets of players $2,\dots,n$ in $\Gamma$ by refinement to ensure perfect recall for $P$: for each original information set $J$ of some player $j\in\{2,\dots,n\}$, we partition $J$ into the equivalence classes induced by the relation $h\sim h^{\prime}$ if and only if the sequence of information sets and actions at which any of the original opponents $\{2,\dots,n\}$ moved along the path to $h$ is identical to that along the path to $h^{\prime}$.
-By construction, $\Gamma^{\prime}$ is a two-player extensive-form game with perfect recall (for players $1$ and $P$), and its size is linear in the size of $\Gamma$.
-Lemma 1.
-Let $\Gamma$ be an $n$-player extensive-form game of imperfect information with perfect recall, and let $\Gamma^{\prime}$ be the two-player game obtained by merging players $2,\dots,n$ into a single opponent $P$ as above.
-Fix an information set $I_{1}$ of player 1 and an action $a\in A(I_{1})$.
-Then $a$ is strictly dominated for player 1 at $I_{1}$ in $\Gamma$ (according to Definition 1) if and only if it is strictly dominated for player 1 at $I_{1}$ in $\Gamma^{\prime}$.
-Proof.
-Let $\Sigma_{j}$ denote the set of behavior strategies of player $j$ in $\Gamma$, and let
-$\Sigma_{-1}\;=\;\Sigma_{2}\times\cdots\times\Sigma_{n}$
-denote the set of behavior-strategy profiles of all players other than 1.
-For an information set $I_{1}$ of player 1, let $\Sigma_{-1}^{I_{1}}\subseteq\Sigma_{-1}$ denote the subset of profiles that “always play to reach $I_{1}$ when possible,” as in Definition 1.
-Similarly, let $\Sigma_{P}$ denote the set of behavior strategies of player $P$ in $\Gamma^{\prime}$, and let $\Sigma_{P}^{I_{1}}\subseteq\Sigma_{P}$ be the subset that always plays to reach $I_{1}$ when possible in $\Gamma^{\prime}$.
-Because behavior strategies assign an independent probability distribution to each information set, and because the information sets of $P$ in $\Gamma^{\prime}$ are obtained by refining and relabeling the information sets of players $2,\dots,n$ in $\Gamma$, there is a natural bijection
-$\Phi:\Sigma_{-1}\;\longrightarrow\;\Sigma_{P}$
-that maps a profile $(\sigma_{2},\dots,\sigma_{n})$ to a single behavior strategy $\sigma_{P}$ which, at each information set of some original player $j\geq 2$, plays according to $\sigma_{j}$.
-The local “always reach $I_{1}$” constraints are preserved under this mapping, so
-$\Phi\big(\Sigma_{-1}^{I_{1}}\big)=\Sigma_{P}^{I_{1}}.$
-For any profile $(\sigma_{1},\sigma_{-1})$ in $\Gamma$, and the corresponding $(\sigma_{1},\sigma_{P})$ in $\Gamma^{\prime}$ with $\sigma_{P}=\Phi(\sigma_{-1})$, the induced distribution over terminal histories is the same in $\Gamma$ and $\Gamma^{\prime}$, since the game tree and player 1’s actions are unchanged and the opponents’ local behavior at each information set is identical.
-Therefore, for all such profiles we have
-$u_{1}(\sigma_{1},\sigma_{-1})\;=\;u_{1}(\sigma_{1},\sigma_{P}),$
-where $u_{1}$ is player 1’s payoff.
-Now fix an action $a\in A(I_{1})$.
-By Definition 1, $a$ is strictly dominated for player 1 at $I_{1}$ in $\Gamma$ if and only if there exists a behavior strategy $\sigma_{1}^{-a}$ for player 1 that always reaches $I_{1}$ and never plays $a$ at $I_{1}$ such that, for every behavior strategy $\sigma_{1}^{a}$ that always reaches $I_{1}$ and plays $a$ with probability 1 at $I_{1}$, and for every $\sigma_{-1}\in\Sigma_{-1}^{I_{1}}$,
-$u_{1}(\sigma_{1}^{-a},\sigma_{-1})\;>\;u_{1}(\sigma_{1}^{a},\sigma_{-1}).$
-Using the bijection $\Phi$ and equality of payoffs, this is equivalent to: there exists $\sigma_{1}^{-a}$ such that for every $\sigma_{1}^{a}$ and for every $\sigma_{P}\in\Sigma_{P}^{I_{1}}$,
-$u_{1}(\sigma_{1}^{-a},\sigma_{P})\;>\;u_{1}(\sigma_{1}^{a},\sigma_{P}),$
-which is precisely the strict-dominance condition for action $a$ at $I_{1}$ in the merged two-player game $\Gamma^{\prime}$ (again using Definition 1 with the opponents collapsed into $P$).
-Thus $a$ is strictly dominated in $\Gamma$ if and only if it is strictly dominated in $\Gamma^{\prime}$. ∎
-Theorem 4.
-There exists a polynomial-time algorithm that determines whether a given action is strictly dominated in an $n$-player perfect-recall extensive-form game of imperfect information.
-Proof.
-Let $\Gamma$ be an $n$-player extensive-form game with perfect recall, let $I_{1}$ be an information set of player 1, and let $a\in A(I_{1})$ be the action to be tested.
-First, construct the merged two-player game $\Gamma^{\prime}$ with players $1$ and $P$ as described above.
-This transformation does not duplicate any nodes of the game tree; it only relabels decision nodes and refines some information sets of the opponents.
-Hence the size of $\Gamma^{\prime}$ (number of nodes, information sets, and actions) is linear in the size of $\Gamma$, and the construction can be performed in polynomial time.
-By Lemma 1, action $a$ is strictly dominated for player 1 at $I_{1}$ in $\Gamma$ (according to Definition 1) if and only if it is strictly dominated for player 1 at $I_{1}$ in $\Gamma^{\prime}$.
-The game $\Gamma^{\prime}$ is a two-player extensive-form game with perfect recall.
-By the two-player result stated earlier, strict dominance of a given action in a two-player perfect-recall extensive-form game of imperfect information can be decided in time polynomial in the size of the game, using the sequence-form linear programs derived in the previous section.
-Therefore, we can decide whether $a$ is strictly dominated in $\Gamma$ by:
-  1. 1.
-    constructing the merged game $\Gamma^{\prime}$ in polynomial time, and
-  2. 2.
-    running the two-player strict-dominance algorithm on $\Gamma^{\prime}$ in polynomial time.
-The total running time is polynomial in the size of $\Gamma$, which proves the theorem. ∎
-Theorem 5.
-There exists a polynomial-time algorithm that determines whether a given action is weakly dominated for player 1 in an $n$-player perfect-recall extensive-form game of imperfect information.
-Proof.
-The weak-dominance test for two-player perfect-recall games is polynomial-time (Theorem 2).
-Weak dominance is preserved under the merging transformation exactly as in Lemma 1: the definition only modifies the payoff inequality from strict to weak, but the induced distributions and opponent strategy sets remain in bijection under the merge.
-Thus an action is weakly dominated in the original $n$-player game if and only if it is weakly dominated in the merged two-player game.
-Applying the two-player weak-dominance test to the merged game gives a polynomial-time algorithm for the $n$-player case.
-∎
-Theorem 6.
-Iterated removal of strictly and weakly dominated actions can be performed in polynomial time in an $n$-player perfect-recall extensive-form game of imperfect information.
-Proof.
-At each iteration of the iterated-removal process in the $n$-player game $\Gamma$, for every information set $I_{1}$ of player 1 and every action $a\in A(I_{1})$, we test whether $a$ is strictly or weakly dominated according to Definitions 1 and 2.
-To perform this test, we construct the merged two-player game $\Gamma^{\prime}$ and apply the corresponding strict or weak-dominance algorithm for two-player games.
-By Theorems 4 and 5, each dominance test in $\Gamma^{\prime}$ runs in time polynomial in the size of $\Gamma$.
-The number of actions in $\Gamma$ is polynomial in the size of the game tree, so in each iteration we perform only polynomially many dominance tests, and each such iteration therefore runs in polynomial time.
-Furthermore, each action can be removed at most once, so the total number of iterations is at most the total number of actions, and hence polynomial in the size of $\Gamma$.
-Finally, Lemma 1 ensures that at every iteration the dominance status of each action is identical in the original $n$-player game $\Gamma$ and in the merged two-player game $\Gamma^{\prime}$.
-Therefore an action is removed in $\Gamma$ at a given iteration if and only if the corresponding action is removed in $\Gamma^{\prime}$.
-Combining these observations, we obtain that the entire iterated-removal process runs in polynomial time and yields the same reduced game whether it is viewed in $\Gamma$ or in $\Gamma^{\prime}$.
-∎
+Iterated removal of strictly and weakly dominated actions can be performed in polynomial time in a two-player perfect-recall extensive-form game of imperfect information with publicly observable actions.
 5 Experiments
@@
 We present a new definition that addresses these limitations.
-We show that both strictly and weakly dominated actions can be identified in polynomial time, and that iterative removal can be performed in polynomial time in $n$-player games.
-Our algorithms identify actions that are dominated by any mixed strategy, not necessarily a pure action.
+In games with perfect recall and publicly observable actions we show that both strictly and weakly dominated actions can be identified in polynomial time, and that iterative removal can be performed in polynomial time in two-player games.
+Our algorithms identify actions that are dominated by any behavioral strategy, not necessarily a pure action.
 We demonstrate that removing dominated actions can play a significant role in reducing the size of realistic imperfect-information games.
@@
 Thus, there is potentially more future ground to explore on efficient game reduction by elimination of poor actions.
+The most immediate open problems for future study are to determine the complexity of finding dominated actions in games in which the assumptions of publicly observable actions and/or perfect recall do not hold, as well as studying games with $n>2$ players.